OpenStudy (anonymous):

at what rate does the sum be twice in 15 years

OpenStudy (anonymous):

which sum?

OpenStudy (anonymous):

thats the question .some sum say x

OpenStudy (phi):

can you post more of the question?

OpenStudy (anonymous):

no full question is this .

OpenStudy (anonymous):

the sum be twice in 15 years.to find rate of interest

OpenStudy (phi):

If you are asking what interest will double the principal after 15 years, we need to know if this is simple interest or compound interest. If it is simple interest (see http://www.mathcaptain.com/algebra/simple-interest.html#simple-interest-formula) the formula is I = P * T *R where I is the amount of interest, P is how much you start with, T is the time R is the rate. to double P, we want I to equal P (so we will have P+I = P+P = 2P after 15 years) P= P*15*R (let T = 15 years) solve for R: 15R=1 R= 1/15 or 0.0666... or about 6.67%